# AI Agentic Vulnerability Injection And Transformation with Optimized Reasoning

### v2

@@ I Introduction
The rapid growth in software complexity, coupled with the sophistication of cyber-attacks, poses a significant threat to the global security and stability of digital infrastructures. In 2024 alone, the total number of publicly reported vulnerabilities rose by 25% [1]. Software vulnerabilities refer to weaknesses in system security requirements, design, implementation, or operation, that could be accidentally triggered or intentionally exploited, resulting in a violation of the system’s security policy [2]. These trends highlight the urgent need for scalable and effective automated vulnerability detection and repair systems. Traditional methods, such as static program analysis, have played an important role. However, their reliance on manually crafted heuristics and the combinatorial complexity of their analysis makes them fundamentally limited, with persistent challenges in scalability, adaptability, and false alarm rates [3, 4].
@@
In response, AI-based approaches for code vulnerability detection [5, 6, 7, 8, 9, 10, 11] and repair [12, 13, 14] have recently attracted considerable attention in the research community. These data-driven techniques, show strong potential in identifying and addressing a wide range of vulnerabilities. However, their effectiveness heavily relies on the quality and scale of the training data [11, 15]. This dependency has become even more critical with the emergence of Large Language Models (LLMs) with billions of parameters to train. The scarcity of high-quality, large-scale datasets therefore remains a central obstacle to advancing automated software security [11, 6, 5]. Existing vulnerability datasets can be grouped into the following categories based on how the code samples were collected and annotated:
@@
  • Synthetic datasets: Both the code and its vulnerability labels are generated artificially, typically using predefined insecure coding patterns. A notable example is the Juliet test suite [16] which is a collection of small synthetic test cases, each demonstrating a specific weakness. It is part of a larger SARD dataset [17] which contains test cases ranging from small synthetic code to production software.
@@
    These datasets offer clean and controllable examples but often lack the structural and semantic complexity of real-world code. As a result, AI models trained on such data overfit to superficial patterns—such as specific variable names or function structures—rather than learning the underlying causes of vulnerabilities [5].
@@
  • Large realistic datasets: Several widely used large-scale vulnerability datasets, including ReVeal [5], BigVul [18], CVEfixes [19], CrossVul [20] and Diversevul [10] rely on heuristic-based labeling functions. These approaches typically involve identifying vulnerability-fixing commits from public security databases, such as the National Vulnerability Database (NVD) [21], and labeling the pre-commit versions of affected functions as vulnerable. While these datasets contain code samples from real-world production software, the reliance on heuristics introduces noise and label inaccuracies. Studies have reported label accuracies as low as 25% for some datasets [11] [22]. Additionally, datasets like D2A [23] and Draper [24] use static analyzers to label code, which results in numerous false positives due to the limitations of static analysis tools. The presence of noisy or incorrect labels in all of these datasets poses challenges for training reliable vulnerability detection models [10] [11].
@@
  • Reliable labels datasets: Some datasets attempt to improve label accuracy through manual expert annotation or by enforcing very strict heuristics. For instance, SVEN [25] manually labels 1,606 functions, which ensures clean labels, but covering only 9 Common Weakness Enumeration (CWE) types. PrimeVul [11] applies rigorous heuristics, claiming human-level labeling accuracy across  7,000 vulnerable and  229,000 benign functions spanning over 140 CWEs. Yet, manual curation is inherently tedious and produces only small datasets, while overly strict heuristics hinder scalability.
@@
  • Vulnerability injection-based datasets: Several approaches for injecting bugs into production software have been proposed and implemented [26, 27, 28, 29]. Methods such as VULGEN [30], VinJ [31] and VGX [15] combine pattern mining with deep learning-based localization to identify injection sites and apply vulnerability-inducing edits. These approaches scale better and produce realistic samples, but they remain constrained in both coverage and reliability. In practice these, injected bugs are limited to single-statement edits and may lack the semantic and contextual richness of real-world vulnerabilities. Moreover, the accuracy of injected datasets remains a concern: for example, VULGEN achieves only a 69% success rate in generating truly vulnerable samples, indicating potential noise and limited reliability.
@@
To address these challenges, we introduce a novel AI agentic workflow designed to automatically inject realistic, diverse, category-specific vulnerabilities for large-scale generation of high-quality vulnerability datasets. This framework, called AVIATOR (AI Agentic Vulnerability Injection And Transformation with Optimized Reasoning) coordinates multiple AI agents - simulating expert reasoning - alongside functional agents and traditional code analysis tools. It was designed by understanding the full process of manual injection and splitting it into elementary subtasks, relying on tools and self-correction loops to provide optimal guidance to LLMs. Its design was validated by three cybersecurity experts (including two co-authors). To further boost performance, it leverages Retrieval-Augmented Generation (RAG) [32] for contextual grounding and employs Low-Rank Adaptation (LoRA) [33] for efficient model fine-tuning. Crucially, AVIATOR requires no pattern mining, freeing it from the rigid constraints of mined edits and allowing scalable generation of vulnerabilities that span more CWE categories, with greater semantic and contextual realism.
@@
To evaluate the performance of AVIATOR in introducing vulnerabilities we conducted an experimental study on code samples from three different benchmarks used in the literature [17, 34, 11]. We leveraged formal verification tool ESBMC [35] for supported vulnerabilities in smaller functions and manual verification by experienced computer security researchers for the rest. We compare AVIATOR with existing automated large-scale realistic software vulnerability dataset creation technique. The results suggest a potential for the AVIATOR to outperform other techniques with regard to accuracy of dataset labels. It achieved between 91% and 95% success rates in injecting vulnerabilities across the different benchmarks.
@@ II-A Properties of High-Quality Vulnerability Datasets
Effective training and rigorous benchmarking of AI-based vulnerability detection and repair methods critically depend on the quality of the underlying datasets. To clearly define the desired dataset characteristics, we introduce the ABCD criteria, similar to [2] and [27]:
@@ II-B Motivation
Creating datasets that rigorously fulfill the ABCD criteria and structural requirements remains an open challenge due to significant limitations inherent to existing approaches [6]. Synthetic datasets, while accurate and easily scalable, lack credibility and offer limited diversity, as they rely on simplistic, predefined vulnerability patterns [5]. Heuristic-based labeling, in contrast, provides credible and diverse samples by mining real-world open-source repositories, but it suffers from a trade-off between accuracy and scalability: stricter heuristics yield smaller but cleaner datasets [11], while relaxed heuristics introduce noise [10]. Static analysis-based labeling can generate large datasets, yet accuracy remains a concern due to high false-positive rates. Manual labeling, though highly accurate, is not scalable for large datasets [36].
Vulnerability injection presents a promising alternative. By design, this data augmentation technique enables scalability through automation and preserves credibility when applied to real-world codebases. However, the effectiveness of this approach depends heavily on the injection technique’s ability to introduce diverse and accurate vulnerabilities. Current state-of-the-art methods remain limited in this regard: recent studies report coverage of only 19 distinct CWE classes and achieve an injection success rate of just 69% [30, 31]. Such limitations constrain both the diversity and reliability of resulting datasets, particularly for benchmarking purposes. Nonetheless, these initial results suggest that with more refined techniques, vulnerability injection has the potential to systematically produce accurate, realistic, diverse, and large-scale datasets that better support the development and benchmarking of modern AI-based vulnerability detection and repair systems, as well as traditional static analyzers.
@@ III-C2 Information Loader (tool)
Examines the benign code to identify elements that prevent the intended vulnerability from manifesting. These include explicit sanitization checks (e.g., bounds validation) as well as broader safety mechanisms such as correct memory management, error handling, or defensive programming patterns. The agent produces a structured inventory of these safeguards, which downstream agents use to determine how the code can be realistically transformed into a vulnerable variant. This process mirrors traditional symbolic static analysis and taint-tracking techniques [37], but applied inversely. Instead of proving correctness, it highlights the logic that enforces safety. This reverse reasoning is critical to selecting feasible and non-trivial injection points.
@@
To specialize the vulnerability injection agent without incurring the computational cost of full model fine-tuning, we apply Low-Rank Adaptation (LoRA) [33]. This technique is particularly effective in settings where the base model is large, but the task-specific dataset is limited in size. Rather than updating the full weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, LoRA learns a low-rank decomposition by injecting two trainable matrices $B\in\mathbb{R}^{d\times r}$ and $A\in\mathbb{R}^{r\times k}$, with $r\ll\min(d,k)$, into the model’s architecture: $W=W_{0}+BA$. This reduces the number of trainable parameters while retaining the benefits of large-scale pretraining.
@@
While SFT optimizes for token-level accuracy, it does not directly focus on structural and semantic nuances of code. Therefore, we also apply reinforcement learning (RL) via Group Relative Policy Optimization (GRPO) [38], using the CodeBLEU [39] metric as a reward function.
@@
with $\alpha+\beta+\gamma+\delta=1$. By default, $\alpha=\beta=\gamma=\delta$ [39].
@@ IV-A Benchmark Construction
We evaluate AVIATOR specifically for C/C++ code, given the inherent susceptibility of these languages to low-level vulnerabilities (e.g., memory errors), focusing on the function level in alignment with prior vulnerability detection and repair studies [40]. We used three complementary popular open-source datasets used in previous studies [41] [34] [11].
  • SARD test suite 100 [17, 41]: Contains 102 compact, synthetic test cases, including 34 vulnerable/benign function pairs that can be detected through formal verification tools such as ESBMC [35]. Their compact and self-contained design, makes them ideal for rapid, fully automated validation and served as our primary development benchmark.
  • FormAI [34]: Includes more realistic and structurally complex synthetic examples. Since it does not provide aligned benign/vulnerable pairs, we manually created benign versions for 37 randomly sampled vulnerable functions, without altering the program’s intent and ensuring no vulnerabilities were detectable by formal verification tools on the resulting benign versions. ESBMC is again used for manual validation, offering a stronger challenge to our injection framework.
  • PrimeVul [11]: Contains over 5,000 real-world C/C++ code pairs annotated with vulnerability labels. Due to the lack of full context (e.g., missing class definitions, global variables, and interprocedural dependencies), as well as bigger code samples, formal verification tools such as ESBMC are not directly applicable. We use the provided: (1) training split of 3,789 pairs across 112 CWEs to fine-tune the Vulnerability Injector, (2) validation split of 480 pairs (63 CWEs) as the retrieval knowledge base in AVIATOR’s RAG setup, and (3) 45 randomly sampled pairs from the test split (24 CWEs) for manual evaluation of injection success.
@@ IV-B Implementation Details
We experimented with two base large language models for the AI agents in our workflow. Qwen2.5-Coder-32B-Instruct [42], is a specialized open-source coding LLM with 32 billion parameters. We additionally experimented with Llama-4-Maverick [43], a general-purpose, mixture-of-experts-based LLM with approximately 400 billion parameters (17 billion active parameters), to explore the impact of model size and specialization on the agents’ reasoning capabilities.
@@
We use gte-Qwen2-1.5B-Instruct as the embedding model for vulnerable/benign code function pairs drawn from the validation split of PrimeVul (§IV-A) that constitutes the RAG knowledge base. At the time of our experiments, this model achieved state-of-the-art scores on the MTEB benchmark for embeddings up to 8K tokens [44]. Since 99% of functions in the PrimeVul training split fall below this length, no chunking was required. We set the retrieval parameter to $k=4$, following the recommended default.
@@ IV-D Evaluation and Metrics
For compatible test cases [17] and [34], we automatically measured vulnerability injection success using the formal verification tool ESBMC v7.5 (Efficient SMT-based Bounded Model Checker), recognized for its effectiveness in detecting vulnerabilities in bounded code contexts [35, 45]. It implements state-of-the-art incremental Bounded Model Checking and $k$-induction proof-rule algorithms. Namely, if the solver reaches termination within a bound less than or equal to $k$, we can definitively prove the absence of software errors. We set the bound to 100 and timeout to 5 minutes.
@@
  • Pass@k [46]: Estimates the probability that at least one of $k$ independently sampled outputs is correct. If $c$ is the number of correct outputs among $k$ samples:
@@
For PrimeVul, where formal verification tools such as ESBMC are not directly applicable due to the lack of full context (e.g., missing class definitions and global variables), as well as bigger code samples, we manually validated a sample of 45 injected functions drawn uniformly at random, following prior work [23, 10, 31, 47]. The evaluation was performed by three authors experienced in C/C++, including two senior security researchers. Each injected sample was categorized as (1) weakness present, (2) no weakness, or (3) undecidable—typically due to reliance on external functions or context beyond the injected function. Since we lack complete test cases for most programs, we assess the presence of weaknesses rather than confirmed vulnerabilities (which require demonstrable exploitability). The reported injection success rate reflects the percentage of samples found to contain a weakness, based only on those that could be conclusively assessed.

### v3

@@ I Introduction
The rapid growth in software complexity, coupled with the sophistication of cyber-attacks, poses a significant threat to the global security and stability of digital infrastructures. In 2024 alone, the total number of publicly reported vulnerabilities rose by 25% [cyberThreatIndex_2024]. Software vulnerabilities refer to weaknesses in system security requirements, design, implementation, or operation, that could be accidentally triggered or intentionally exploited, resulting in a violation of the system’s security policy [sate6_delaitre2023]. These trends highlight the urgent need for scalable and effective automated vulnerability detection and repair systems. Traditional methods, such as static program analysis, have played an important role. However, their reliance on manually crafted heuristics and the combinatorial complexity of their analysis makes them fundamentally limited, with persistent challenges in scalability, adaptability, and false alarm rates [Not_using_Static_tools, CastleBenchmarking_dubniczky2025].
@@
In response, AI-based approaches for code vulnerability detection [DLvulnDetectStudy_chakraborty2020, llmVulnSurvey_zhou2024, VulDeePecker_Li2018, SySeVR_Li2022, Rahman_2024, DiverseVul_yizheng2023, PrimeVul_ding2024] and repair [deepcodeaifixfixing_berabi2024, vulnfixvisiontransformer_fu2025, codesecurityvulnerabilityrepair_islam2024] have recently attracted considerable attention in the research community. These data-driven techniques, show strong potential in identifying and addressing a wide range of vulnerabilities. However, their effectiveness heavily relies on the quality and scale of the training data [PrimeVul_ding2024, vgx_nong2024]. This dependency has become even more critical with the emergence of Large Language Models (LLMs) with billions of parameters to train. The scarcity of high-quality, large-scale datasets therefore remains a central obstacle to advancing automated software security [PrimeVul_ding2024, llmVulnSurvey_zhou2024, DLvulnDetectStudy_chakraborty2020]. Existing vulnerability datasets can be grouped into the following categories based on how the code samples were collected and annotated:
@@
  • Synthetic datasets: Both the code and its vulnerability labels are generated artificially, typically using predefined insecure coding patterns. A notable example is the Juliet test suite [juliet2017] which is a collection of small synthetic test cases, each demonstrating a specific weakness. It is part of a larger SARD dataset [nist_sard_2018] which contains test cases ranging from small synthetic code to production software.
@@
    These datasets offer clean and controllable examples but often lack the structural and semantic complexity of real-world code. As a result, AI models trained on such data overfit to superficial patterns—such as specific variable names or function structures—rather than learning the underlying causes of vulnerabilities [DLvulnDetectStudy_chakraborty2020].
@@
  • Large realistic datasets: Several widely used large-scale vulnerability datasets, including ReVeal [DLvulnDetectStudy_chakraborty2020], BigVul [BigVul_jiahao2020], CVEfixes [CVEfixes_guru2021], CrossVul [CrossVul_georgios2021] and Diversevul [DiverseVul_yizheng2023] rely on heuristic-based labeling functions. These approaches typically involve identifying vulnerability-fixing commits from public security databases, such as the National Vulnerability Database (NVD) [nist_nvd_2022], and labeling the pre-commit versions of affected functions as vulnerable. While these datasets contain code samples from real-world production software, the reliance on heuristics introduces noise and label inaccuracies. Studies have reported label accuracies as low as 25% for some datasets [PrimeVul_ding2024] [DataQuality_croft2023]. Additionally, datasets like D2A [d2adatasetbuilt_zheng2021] and Draper [draper_russell2018] use static analyzers to label code, which results in numerous false positives due to the limitations of static analysis tools. The presence of noisy or incorrect labels in all of these datasets poses challenges for training reliable vulnerability detection models [DiverseVul_yizheng2023] [PrimeVul_ding2024].
@@
  • Reliable labels datasets: Some datasets attempt to improve label accuracy through manual expert annotation or by enforcing very strict heuristics. For instance, SVEN [SVEN_he2023] manually labels 1,606 functions, which ensures clean labels, but covering only 9 Common Weakness Enumeration (CWE) types. PrimeVul [PrimeVul_ding2024] applies rigorous heuristics, claiming human-level labeling accuracy across  7,000 vulnerable and  229,000 benign functions spanning over 140 CWEs. Yet, manual curation is inherently tedious and produces only small datasets, while overly strict heuristics hinder scalability.
@@
  • Vulnerability injection-based datasets: Several approaches for injecting bugs into production software have been proposed and implemented [pewny2016evilcoder, dolan2016lava, roy2018bug, kashyap2019automated]. Methods such as VULGEN [vulgen_nong2023], VinJ [vinj_nong2024] and VGX [vgx_nong2024] combine pattern mining with deep learning-based localization to identify injection sites and apply vulnerability-inducing edits. These approaches scale better and produce realistic samples, but they remain constrained in both coverage and reliability. In practice these, injected bugs are limited to single-statement edits and may lack the semantic and contextual richness of real-world vulnerabilities. Moreover, the accuracy of injected datasets remains a concern: for example, VULGEN achieves only a 69% success rate in generating truly vulnerable samples, indicating potential noise and limited reliability.
@@
To address these challenges, we introduce a novel AI agentic workflow designed to automatically inject realistic, diverse, category-specific vulnerabilities for large-scale generation of high-quality vulnerability datasets. This framework, called AVIATOR (AI Agentic Vulnerability Injection And Transformation with Optimized Reasoning) coordinates multiple AI agents - simulating expert reasoning - alongside functional agents and traditional code analysis tools. It was designed by understanding the full process of manual injection and splitting it into elementary subtasks, relying on tools and self-correction loops to provide optimal guidance to LLMs. Its design was validated by three cybersecurity experts (including two co-authors). To further boost performance, it leverages Retrieval-Augmented Generation (RAG) [rag_lewis2021] for contextual grounding and employs Low-Rank Adaptation (LoRA) [lora_hu2021] for efficient model fine-tuning. Crucially, AVIATOR requires no pattern mining, freeing it from the rigid constraints of mined edits and allowing scalable generation of vulnerabilities that span more CWE categories, with greater semantic and contextual realism.
@@
To evaluate the performance of AVIATOR in introducing vulnerabilities we conducted an experimental study on code samples from three different benchmarks used in the literature [nist_sard_2018, formai_tihanyi2023, PrimeVul_ding2024]. We leveraged formal verification tool ESBMC [esbmc_gadelha2018] for supported vulnerabilities in smaller functions and manual verification by experienced computer security researchers for the rest. We compare AVIATOR with existing automated large-scale realistic software vulnerability dataset creation technique. The results suggest a potential for the AVIATOR to outperform other techniques with regard to accuracy of dataset labels. It achieved between 91% and 95% success rates in injecting vulnerabilities across the different benchmarks.
@@ II-A Properties of High-Quality Vulnerability Datasets
Effective training and rigorous benchmarking of AI-based vulnerability detection and repair methods critically depend on the quality of the underlying datasets. To clearly define the desired dataset characteristics, we introduce the ABCD criteria, similar to [sate6_delaitre2023] and [dolan2016lava]:
@@ II-B Motivation
Creating datasets that rigorously fulfill the ABCD criteria and structural requirements remains an open challenge due to significant limitations inherent to existing approaches [llmVulnSurvey_zhou2024]. Synthetic datasets, while accurate and easily scalable, lack credibility and offer limited diversity, as they rely on simplistic, predefined vulnerability patterns [DLvulnDetectStudy_chakraborty2020]. Heuristic-based labeling, in contrast, provides credible and diverse samples by mining real-world open-source repositories, but it suffers from a trade-off between accuracy and scalability: stricter heuristics yield smaller but cleaner datasets [PrimeVul_ding2024], while relaxed heuristics introduce noise [DiverseVul_yizheng2023]. Static analysis-based labeling can generate large datasets, yet accuracy remains a concern due to high false-positive rates. Manual labeling, though highly accurate, is not scalable for large datasets [automaticdatalabelingsoftware_le2024].
Vulnerability injection presents a promising alternative. By design, this data augmentation technique enables scalability through automation and preserves credibility when applied to real-world codebases. However, the effectiveness of this approach depends heavily on the injection technique’s ability to introduce diverse and accurate vulnerabilities. Current state-of-the-art methods remain limited in this regard: recent studies report coverage of only 19 distinct CWE classes and achieve an injection success rate of just 69% [vulgen_nong2023, vinj_nong2024]. Such limitations constrain both the diversity and reliability of resulting datasets, particularly for benchmarking purposes. Nonetheless, these initial results suggest that with more refined techniques, vulnerability injection has the potential to systematically produce accurate, realistic, diverse, and large-scale datasets that better support the development and benchmarking of modern AI-based vulnerability detection and repair systems, as well as traditional static analyzers.
@@ III-C2 Information Loader (tool)
Examines the benign code to identify elements that prevent the intended vulnerability from manifesting. These include explicit sanitization checks (e.g., bounds validation) as well as broader safety mechanisms such as correct memory management, error handling, or defensive programming patterns. The agent produces a structured inventory of these safeguards, which downstream agents use to determine how the code can be realistically transformed into a vulnerable variant. This process mirrors traditional symbolic static analysis and taint-tracking techniques [staticAnalysis_schwartz2010], but applied inversely. Instead of proving correctness, it highlights the logic that enforces safety. This reverse reasoning is critical to selecting feasible and non-trivial injection points.
@@
To specialize the vulnerability injection agent without incurring the computational cost of full model fine-tuning, we apply Low-Rank Adaptation (LoRA) [lora_hu2021]. This technique is particularly effective in settings where the base model is large, but the task-specific dataset is limited in size. Rather than updating the full weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, LoRA learns a low-rank decomposition by injecting two trainable matrices $B\in\mathbb{R}^{d\times r}$ and $A\in\mathbb{R}^{r\times k}$, with $r\ll\min(d,k)$, into the model’s architecture: $W=W_{0}+BA$. This reduces the number of trainable parameters while retaining the benefits of large-scale pretraining.
@@
While SFT optimizes for token-level accuracy, it does not directly focus on structural and semantic nuances of code. Therefore, we also apply reinforcement learning (RL) via Group Relative Policy Optimization (GRPO) [grpo_shao2024], using the CodeBLEU [codebleu_ren2020] metric as a reward function.
@@
with $\alpha+\beta+\gamma+\delta=1$. By default, $\alpha=\beta=\gamma=\delta$ [codebleu_ren2020].
@@ IV-A Benchmark Construction
We evaluate AVIATOR specifically for C/C++ code, given the inherent susceptibility of these languages to low-level vulnerabilities (e.g., memory errors), focusing on the function level in alignment with prior vulnerability detection and repair studies [CppUnsafe_erlingsson2025]. We used three complementary popular open-source datasets used in previous studies [sard100_hoole2016] [formai_tihanyi2023] [PrimeVul_ding2024].
  • SARD test suite 100 [nist_sard_2018, sard100_hoole2016]: Contains 102 compact, synthetic test cases, including 34 vulnerable/benign function pairs that can be detected through formal verification tools such as ESBMC [esbmc_gadelha2018]. Their compact and self-contained design, makes them ideal for rapid, fully automated validation and served as our primary development benchmark.
  • FormAI [formai_tihanyi2023]: Includes more realistic and structurally complex synthetic examples. Since it does not provide aligned benign/vulnerable pairs, we manually created benign versions for 37 randomly sampled vulnerable functions, without altering the program’s intent and ensuring no vulnerabilities were detectable by formal verification tools on the resulting benign versions. ESBMC is again used for manual validation, offering a stronger challenge to our injection framework.
  • PrimeVul [PrimeVul_ding2024]: Contains over 5,000 real-world C/C++ code pairs annotated with vulnerability labels. Due to the lack of full context (e.g., missing class definitions, global variables, and interprocedural dependencies), as well as bigger code samples, formal verification tools such as ESBMC are not directly applicable. We use the provided: (1) training split of 3,789 pairs across 112 CWEs to fine-tune the Vulnerability Injector, (2) validation split of 480 pairs (63 CWEs) as the retrieval knowledge base in AVIATOR’s RAG setup, and (3) 45 randomly sampled pairs from the test split (24 CWEs) for manual evaluation of injection success.
@@ IV-B Implementation Details
We experimented with two base large language models for the AI agents in our workflow. Qwen2.5-Coder-32B-Instruct [qwen25coder_hui2024], is a specialized open-source coding LLM with 32 billion parameters. We additionally experimented with Llama-4-Maverick [llama4_meta2025], a general-purpose, mixture-of-experts-based LLM with approximately 400 billion parameters (17 billion active parameters), to explore the impact of model size and specialization on the agents’ reasoning capabilities.
@@
We use gte-Qwen2-1.5B-Instruct as the embedding model for vulnerable/benign code function pairs drawn from the validation split of PrimeVul (§IV-A) that constitutes the RAG knowledge base. At the time of our experiments, this model achieved state-of-the-art scores on the MTEB benchmark for embeddings up to 8K tokens [mteb_huggingface2023]. Since 99% of functions in the PrimeVul training split fall below this length, no chunking was required. We set the retrieval parameter to $k=4$, following the recommended default.
@@ IV-D Evaluation and Metrics
For compatible test cases [nist_sard_2018] and [formai_tihanyi2023], we automatically measured vulnerability injection success using the formal verification tool ESBMC v7.5 (Efficient SMT-based Bounded Model Checker), recognized for its effectiveness in detecting vulnerabilities in bounded code contexts [esbmc_gadelha2018, Sw_verification_competion_beyer2023]. It implements state-of-the-art incremental Bounded Model Checking and $k$-induction proof-rule algorithms. Namely, if the solver reaches termination within a bound less than or equal to $k$, we can definitively prove the absence of software errors. We set the bound to 100 and timeout to 5 minutes.
@@
  • Pass@k [passk_chen2021]: Estimates the probability that at least one of $k$ independently sampled outputs is correct. If $c$ is the number of correct outputs among $k$ samples:
@@
For PrimeVul, where formal verification tools such as ESBMC are not directly applicable due to the lack of full context (e.g., missing class definitions and global variables), as well as bigger code samples, we manually validated a sample of 45 injected functions drawn uniformly at random, following prior work [d2adatasetbuilt_zheng2021, DiverseVul_yizheng2023, vinj_nong2024, lessIsEnough_wen2023]. The evaluation was performed by three authors experienced in C/C++, including two senior security researchers. Each injected sample was categorized as (1) weakness present, (2) no weakness, or (3) undecidable—typically due to reliance on external functions or context beyond the injected function. Since we lack complete test cases for most programs, we assess the presence of weaknesses rather than confirmed vulnerabilities (which require demonstrable exploitability). The reported injection success rate reflects the percentage of samples found to contain a weakness, based only on those that could be conclusively assessed.